Two cities are 3040 miles apart. A plane leaves one of them,traveling towards the other at an average speed of 350 miles per hour. At the same time a plane leaves the other, traveling towards the first, at an average speed of 410 miles per hour. How long will it take them to meet ?
Answer by stanbon(75887)

Two cities are 3040 miles apart. A plane leaves one of them,traveling towards the other at an average speed of 350 miles per hour. At the same time a plane leaves the other, traveling towards the first, at an average speed of 410 miles per hour. How long will it take them to meet ?
--------
closing rate = 350+410 = 760 mph
----
time to meet = distance/rate = 3040/760 = 4 hrs.
Cheers,
Stan H.